Consider the titration of a 28.0 −mL sample of 0.170 MCH3NH2 with 0.150 M HBr. Kb for methylamine, CH3NH2, at 25 C is 4.4 x 10-4 Determine each of the following.

the pH at 6.0 mL of added acid

the pH at one-half of the equivalence point

the pH at the equivalence point

the pH after adding 6.0 mL of acid beyond the equivalence point
Answer #1

1)

mmoles of CH3NH2 = 28 x 0.170 = 4.76

mmoles of HBr = 6 x 0.150 = 0.90

pKb = 3.356

pKa = 10.644

CH3NH2 +   HBr    ------------> CH3NH3+

4.76             0.90                          0

3.86               0                            0.90

pH = pKa + log [base / acid]

    = 10.644 + log [3.86 / 0.90]

pH = 11.28

2)

At half - equivalence point :

pH = pKa

pH = 10.64

3)

At equivalence point :

volume of acid = 31.73 mL

here only salt remains .

salt concentration = 4.76 / 31.73 + 28 = 0.0797 M

pH = 7 - 1/2 (pKb + log C)

    = 7 - 1/2 (3.356 + log 0.0797)

pH = 5.87

4)

mmoles of acid = 37.73 x 0.150 = 5.66

[H+] = 5.66 - 4.76 / 37.73 + 28 = 0.0137 M

pH = -log (0.0137)

pH = 1.86